OpenStudy (anonymous):

What is the approximate volume of the cylinder? height = 6.4 m; diameter = 4.6 m A. 34 m3 B. 46 m3 C. 92 m3 D. 106 m3

OpenStudy (anonymous):

Volume = A * H = pi(radius)^2 * height

OpenStudy (anonymous):

i think its d ask @SiennaSurfer hes great at math

OpenStudy (anonymous):

Volume=pi*(2.3)^2*6.4

OpenStudy (anonymous):

is approximately 106.36176088

OpenStudy (anonymous):

so d.

OpenStudy (anonymous):

cool i was right

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 9 cm. What is the approximate radius of the cylinder? A. 4.5 cm B. 5.1 cm C. 12.9 cm D. 25.8 cm